# Writing – original draft, Data curation, Formal analysis, Methodology, Conceptualization

### v1

@@ 2.4 Model Selection
All models were implemented using the Scikit-learn (sklearn) library. These models were chosen to reflect a diverse range of learning paradigms and to enable comparative analysis of their predictive capabilities on the EHG dataset.
2.5 Evaluation Metrics
To evaluate the performance of each binary classifier, we calculated the Classification Accuracy, Precision, Recall, F1-score, and the Area Under the Curve (AUC). The evaluation was based on the counts of correctly predicted premature records (TP), correctly predicted term records (TN), Term records incorrectly predicted as premature (FP), and Premature records incorrectly predicted as term (FN). The metrics were calculated using the following formulas:
@@ 2.5 Evaluation Metrics
$\text{AUC}=\int_{-\infty}^{\infty}\text{TP}(T)\cdot\text{FP}^{\prime}(T)\,dT$ (19)
The AUC represents the probability that a classifier will rank a randomly chosen positive instance higher than a randomly chosen negative one, showing the discriminating power between two classes. Equivalently, the AUC equals the area under the Receiver Operating Characteristic (ROC) curve, representing the classifier’s ability to discriminate between positive and negative classes.
@@
To address class imbalance in TPEHGT dataset, we implemented an iterative balanced sampling strategy:
@@
  • Minority class inclusion: All available preterm samples were used (94 total: 47 contraction intervals + 47 dummy intervals)
@@
  • Majority class sampling: An equal number of term samples (94) were randomly selected from the larger pool (106 total: 53 contraction + 53 dummy intervals) to create balanced subsets
To ensure fairness and reduce sampling bias, this process was repeated over 20 independent iterations, each using a different random subset of the majority class. The final evaluation metrics were reported as the average across these iterations. Given the limited dataset size, we applied five fold stratified cross validation within each iteration to preserve class distribution across folds. A fixed random seed was used to ensure reproducibility and enable direct comparison across different classifiers. This approach ensures statistical robustness while mitigating sampling bias in performance evaluation.
3 Results and Discussion
@@
This section presents the classification results and key findings from our experiments. First, Section 3.1 reports baseline performance on the TPEHGT dataset, including results with and without KLT based denoising. Section 3.2 provides ablation studies to analyze the contribution of different factors. In particular, Section 3.2.1 quantifies the performance gain from applying KLT, after which all subsequent ablation results are reported using KLT transformed signals only. Section 3.2.2 and Section 3.2.3 then evaluate the influence of fixed interval segmentation with and without TOCO signal, respectively. Next, Section 3.3 assesses the generalization of the proposed approach on the TPEHG database. Finally, Section 3.4 compares our best performing model against existing methods in the literature.
3.1 Performance Evaluation on TPEHGT dataset
@@
The classification methods discussed earlier were implemented using the TPEHGT dataset [12]. Since this dataset includes labeled contraction intervals, the signals were segmented into contraction and dummy intervals accordingly. This segmentation resulted in a balanced dataset of 188 samples in total, comprising 94 term and 94 preterm intervals. Feature extraction was carried out using 20 MFCCs, along with six statistics of db8 wavelet coefficients extracted from five detail sub bands (D1–D5) and one approximation sub band (A5) per channel, resulting in 36 wavelet based features (6 statistics $\times$ 6 sub bands), plus the PA of the normalized power spectrum. Combined with 20 MFCCs, this yields a total of 57 features per channel. Considering all EHG channels together, the total number of features across the dataset amounts to 57 $\times$ [number of channels].
@@
To evaluate classification performance, we tested eight classical ML models. Robustness was ensured through a five fold stratified cross validation with balanced sampling, repeated 20 times. Model performance was compared on both raw filtered signals and their KLT versions, which denoise and decorrelate the signals to improve feature quality. The results are summarized in Table 1.
@@
Table 1: Classification performance of eight models on the 2018 TPEHGT dataset using contraction-dummy interval sampling. Results are reported for both the original signals and those processed with the KLT.
@@
The results shown in Table 1 reflect model performance when EHG signals were segmented based on contraction and non contraction(dummy) intervals, which simulate annotated contraction periods and the intervening rest periods. This segmentation approach inherently prioritizes physiologically relevant signal segments, which likely explains the overall strong performance across models.
@@
3.2 Ablation Studies
3.2.1 The impact of KLT
In the proposed pipeline, denoising was performed using the KLT to suppress residual noise and overlapping signal components that persist after band pass filtering. To assess its effectiveness, we evaluated classification performance both with and without KLT, as shown in Table 1.
@@ 3.2.1 The impact of KLT
3.2.2 Performance with TOCO signals
In real time PTB detection, identifying contraction intervals from EHG signals using TOCO is not always practical. While some studies have attempted to detect contractions directly from EHG signals [10], this remains a challenging task. To evaluate the potential benefits of including TOCO, we analyzed EHG recordings segmented blindly into non overlapping 3-minute intervals, simulating a clinical monitoring setup. This resulted in a balanced dataset of 260 samples (130 term + 130 preterm). Both EHG channels and the TOCO signal were utilized during the feature extraction process. Feature extraction followed the same pipeline as in the contraction based setup, using 20 MFCCs, six statistical features from db8 wavelet coefficients across five detail sub bands (D1–D5) and one approximation sub-band (A5), and the PA of the normalized power spectrum, resulting in a total of 57 features per channel.
@@ 3.2.2 Performance with TOCO signals
Table 2 presents the classification results for the TPEHGT dataset using fixed 3-minute intervals, where both EHG and TOCO signals were included during feature extraction.
Table 2: Classification performance of eight models on the 2018 TPEHGT dataset (with TOCO signal) using fixed 3-minute interval sampling. Results are shown for those transformed with the KLT. Evaluation metrics include Accuracy, Precision, Recall, F1 Score, and AUC.
@@
Comparing Tables 1 and 2, we observe a trade off between contraction based and fixed interval segmentation: some models excel in one setup but not the other. It is important to note that direct comparison is not meaningful because the datasets differ in both sample size and segmentation strategy. The contraction based dataset contains 188 samples, while the fixed size window dataset includes 260 samples. Despite these differences, the aim is not to rank the segmentation strategies but to highlight the overall versatility of the proposed approach. The combination of KLT based denoising, which compensates for noise vulnerability in specific models, and contextual feature selection, which emphasizes physiologically relevant patterns, ensures that model performance remains consistently high across different scenarios. This demonstrates the method’s adaptability to diverse real world conditions, ranging from fully monitored clinical settings to real time or low resource applications.
3.2.3 Performance without TOCO signals
In real time PTB detection, access to TOCO signals may not always be feasible, particularly in low resource or wearable monitoring setups. To simulate such scenarios, we evaluated the classification performance using only EHG channels, while keeping the fixed 3-minute interval segmentation consistent with the previous setup. This approach avoids reliance on external contraction measurements and tests the model’s robustness in practical, TOCO independent conditions.
Table 3 summarizes the classification results for the TPEHGT dataset without TOCO signals. Overall, the models achieved accuracies ranging from 88% to 96%, which is only slightly lower than the with TOCO setup (Table 2, 87–97%). Similar to the with TOCO case, ensemble methods (RF, GB, CB) consistently outperformed classical models, with GB achieving the highest accuracy of 96.82% and a strong AUC of 0.9953 on the KLT denoised signals.
Table 3: Classification performance of eight models on the 2018 TPEHGT dataset (without TOCO signal) using fixed 3-minute interval sampling. Results are shown for those transformed with the KLT. Evaluation metrics include Accuracy, Precision, Recall, F1 Score, and AUC.
@@ 3.2.3 Performance without TOCO signals
3.3 Performance Validation on the TPEHG Database
To validate the robustness of the proposed feature selection method, we utilized the 2012 TPEHG dataset. Since this dataset lacks TOCO signals, we applied the same fixed duration segmentation approach described in Section 3.2.3, utilizing 3-minute non overlapping windows. The original dataset contained 300, thirty minute recordings (38 preterm and 262 term). Through our balanced sampling strategy, we generated 720 segments (380 preterm and 380 term) per training iteration. From each segmented signal, we extracted 57 features per channel as previously described. The corresponding performance metrics are presented in Table 4.
Table 4: Classification results for eight models applied to the 2012 TPEHG dataset using fixed 3-minute interval sampling. KLT processed signal variants are evaluated across Accuracy, Precision, Recall, F1 Score, and AUC metrics.
@@ 3.3 Performance Validation on the TPEHG Database
These results also highlight the limitations of the available datasets and the challenges posed by sampling strategies. Despite balancing the classes during training, the original imbalance in subject distribution, particularly the low number of preterm cases, can still influence model generalization. This limitation is further emphasized in Section 3.4, where we compare our approach to existing models. The analysis reveals that improper oversampling strategies employed in prior works have significantly inflated reported performance, often resulting in overly inflated accuracy metrics that do not accurately reflect real world deployment conditions.
@@
3.4 Comparison with existing models
Building upon the methodological concerns raised earlier regarding inflated performance metrics from improper oversampling, our comparative analysis adopts a rigorous approach to benchmark evaluation. We restrict our comparison to studies that have either been methodologically validated by Vandewiele et al. [14] or that explicitly document appropriate sampling strategies in their original publications. Table 5 presents this curated comparison, including both originally reported results and, where available, the methodologically corrected results from Vandewiele et al. In addition, we incorporate recent classical ML studies that report reproducible methodologies and clinically interpretable outcomes. In addition to the classifiers explored in this study, results from prior works using AdaBoost (AB) are also included for completeness.
@@ 3.4 Comparison with existing models
Table 5: Comparison of reported and tuned accuracy from previous studies with evaluation metric.
Study Classifier Metric Uncorrected Results (%) Corrected Results (%)
[24] SVM AUC 92.00 60.75
[25] AB AUC 96.60 57.54
[26] RF AUC 94.00 52.31
[27] SVM AUC 99.00 56.04
[28] QDA AUC 99.44 63.86
[29] CB AUC NA 70.00
[30] Bayes AUC NA 84.00
[16] SVM AUC NA 86.00
[31] SVM AUC NA 89.00
[32] QDA AUC NA 89.00
[33] HVG + SVM AUC NA 97.00
Proposed CB AUC NA 99.88
PTB detection is a critical and sensitive clinical task; yet, much of the prior literature has reported overly optimistic results, portraying PTB prediction as a solved problem. In reality, this is far from the truth. As shown in Table 5, many earlier studies relied on flawed sampling strategies, particularly improper oversampling and data leakage, leading to dramatically inflated performance metrics. For instance, several reported AUCs exceeding 90% dropped on average by more than 25 percentage points, in some cases falling below 65% when evaluated using proper methodology. This striking discrepancy highlights the extent of overestimation in prior work and underscores the urgent need to reopen and rigorously reevaluate the PTB prediction problem. Notably, the limitations of current datasets continue to be a significant barrier to the development of generalizable and reliable models. This study represents an essential first step in that re-evaluation, while also calling for more extensive datasets and methodological rigor to add credibility to AI based PTB research. In doing so, it aims to build clinical trust by prioritizing explainability, transparency, and interoperability in model design.
@@
4 Conclusion
Preterm birth (PTB) affects approximately 10% of live births and remains the leading cause of death among children under five. Survivors often face lifelong challenges, including developmental delays, chronic diseases, and sensory impairments. Early and accurate prediction of PTB is therefore essential for improving neonatal outcomes and guiding timely interventions. This study proposed a robust and reproducible machine learning pipeline for PTB prediction using EHG signals, addressing key limitations in prior research, particularly issues with improper oversampling strategies and limited focus on frequency domain features.
@@ 4 Conclusion
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
[…]

### v2

@@ II-E Model Selection
All models were implemented using the Scikit learn (sklearn) library. These models were chosen to reflect a diverse range of learning paradigms and to enable comparative analysis of their predictive capabilities on the EHG dataset.
II-F Evaluation Protocol and Metrics
Segmentation defines the sample unit as described in Section II-C. All reported results are computed at the segment level (contraction/dummy intervals or fixed 3-minute windows), where each segment inherits the corresponding record label (term vs. preterm).
To evaluate each binary classifier, we report Classification Accuracy, Precision, Recall, F1-score, and the Area Under the Curve (AUC). Let $TP$, TNTN, FPFP, and $FN$ denote the counts of correctly predicted preterm segments, correctly predicted term segments, term segments misclassified as preterm, and preterm segments misclassified as term, respectively. The metrics are computed as:
@@ II-F Evaluation Protocol and Metrics
$\text{AUC}=\int_{0}^{1}\text{TPR}(\text{FPR})\,d(\text{FPR})$ (21)
AUC summarizes the discriminative ability of the classifier across all decision thresholds and equals the area under the Receiver Operating Characteristic (ROC) curve.
@@
Iterative balanced sampling: To address class imbalance for the TPEHGT contraction–dummy interval setting (Section II-C1), we employ an iterative balanced sampling strategy:
@@
  • Minority class: all available preterm segments are used (94 total: 47 contraction + 47 dummy).
@@
  • Majority class: an equal number of term segments (94) are randomly selected from the available pool (106 total: 53 contraction + 53 dummy).
This balancing is repeated for twenty independent iterations using different random seeds for the term undersampling. Reported results are averaged across the twenty iterations.
Cross validation: Within each balanced subset, we perform five fold stratified cross validation at the segment level to preserve the class distribution across folds. The cross-validation split is generated using a fixed random seed for reproducibility, while the majority class subsampling varies across iterations.
III Results and Discussion
@@
This section presents the classification results and key findings from our experiments. First, Section III-A reports baseline performance on the TPEHGT dataset, including results with and without KLT based denoising. Section III-B provides ablation studies to analyze the contribution of different factors. In particular, Section III-B1 quantifies the performance gain from applying KLT, after which all subsequent ablation results are reported using KLT transformed signals only. Section III-B2 and Section III-B3 then evaluate the influence of fixed interval segmentation with and without TOCO signal, respectively. Next, Section III-C assesses the generalization of the proposed approach on the TPEHG database. Finally, Section III-D compares our best performing model against existing methods in the literature.
III-A Performance Evaluation on TPEHGT dataset
@@
The classification methods discussed earlier were implemented using the TPEHGT dataset [12]. We report results using the contraction-dummy interval samples defined in Sec. II-C1, yielding 188 segment samples in total. Feature extraction was carried out using 20 MFCCs, along with six statistics of db8 wavelet coefficients extracted from five detail sub bands (D1–D5) and one approximation sub band (A5) per channel, resulting in 36 wavelet based features (6 statistics $\times$ 6 sub bands), plus the PA of the normalized power spectrum. Combined with 20 MFCCs, this yields a total of 57 features per channel. Considering three EHG channels, the total feature dimension is $57\times 3=171$.
@@
To evaluate classification performance, we tested eight classical ML models. Performance is reported using the evaluation protocol in Section II-F, including iterative balanced sampling and five fold stratified cross validation. Model performance was compared on both raw filtered signals and their KLT versions, which denoise and decorrelate the signals to improve feature quality. The results are summarized in Table I.
@@
TABLE I: Classification performance of eight models on the 2018 TPEHGT dataset using contraction-dummy interval sampling. Results are reported for both the original signals and those processed with the KLT.
@@
The results shown in Table I reflect model performance when EHG signals were segmented based on contraction and non contraction (dummy) intervals, which simulate annotated contraction periods and the intervening rest periods. This segmentation approach inherently prioritizes physiologically relevant signal segments, which likely explains the overall strong performance across models.
@@
III-B Ablation Studies
III-B1 The impact of KLT
In the proposed pipeline, denoising was performed using the KLT to suppress residual noise and overlapping signal components that persist after band pass filtering. To assess its effectiveness, we evaluated classification performance both with and without KLT, as shown in Table I.
@@ III-B1 The impact of KLT
III-B2 Performance with TOCO signals
In real time PTB detection, identifying contraction intervals from EHG signals using TOCO is not always practical. While some studies have attempted to detect contractions directly from EHG signals [10], this remains a challenging task. To evaluate the potential benefits of including TOCO, we adopt the fixed 3-minute non overlapping segmentation described in Section II-C2, which simulates a blind real time monitoring setup. For this setting, both EHG channels and the TOCO signal were used during feature extraction. For TPEHGT, this produces 260 segment samples (130 term + 130 preterm) under the fixed window setup. Feature extraction followed the same pipeline as in the contraction based setup, using twenty MFCCs, six statistical features from db8 wavelet coefficients across five detail sub bands (D1–D5) and one approximation sub-band (A5), and the PA of the normalized power spectrum, resulting in a total of 57 features per channel.
@@ III-B2 Performance with TOCO signals
Table II presents the classification results for the TPEHGT dataset using fixed 3-minute intervals, where both EHG and TOCO signals were included during feature extraction.
TABLE II: Classification performance of eight models on the 2018 TPEHGT dataset (with TOCO signal) using fixed 3-minute interval sampling. Results are shown for those transformed with the KLT. Evaluation metrics include Accuracy, Precision, Recall, F1 Score, and AUC.
@@
Comparing Tables I and II, we observe a trade off between contraction based and fixed interval segmentation: some models excel in one setup but not the other. It is important to note that direct comparison is not meaningful because the datasets differ in both sample size and segmentation strategy. The contraction based dataset contains 188 samples, while the fixed size window dataset includes 260 samples. Despite these differences, the aim is not to rank the segmentation strategies but to highlight the overall versatility of the proposed approach. The combination of KLT based denoising, which compensates for noise vulnerability in specific models, and contextual feature selection, which emphasizes physiologically relevant patterns, ensures that model performance remains consistently high across different scenarios. This demonstrates the method’s adaptability to diverse real world conditions, ranging from fully monitored clinical settings to real time or low resource applications.
III-B3 Performance without TOCO signals
In real time PTB detection, access to TOCO signals may not always be feasible, particularly in low resource or wearable monitoring setups. To simulate such scenarios, we use the same fixed 3-minute segmentation described in Section II-C2 while excluding TOCO features and using only the EHG channels. This approach avoids reliance on external contraction measurements and tests the model’s robustness in practical, TOCO independent conditions.
Table III summarizes the classification results for the TPEHGT dataset without TOCO signals. Overall, the models achieved accuracies ranging from 88% to 96%, which is only slightly lower than the with TOCO setup (Table II, 87–97%). Similar to the with TOCO case, ensemble methods (RF, GB, CB) consistently outperformed classical models, with GB achieving the highest accuracy of 96.82% and a strong AUC of 0.9953 on the KLT denoised signals.
TABLE III: Classification performance of eight models on the 2018 TPEHGT dataset (without TOCO signal) using fixed 3-minute interval sampling. Results are shown for those transformed with the KLT. Evaluation metrics include Accuracy, Precision, Recall, F1 Score, and AUC.
@@ III-B3 Performance without TOCO signals
III-C EVALUATION ON THE TPEHG DATASET
To further evaluate the proposed pipeline on an additional dataset, we used the 2012 TPEHG dataset. Since this dataset lacks TOCO signals, we apply the fixed 3-minute non overlapping segmentation described in Section II-C2 and extract features from EHG channels only. The original dataset contained 300, thirty minute recordings (38 preterm and 262 term). Through our balanced sampling strategy, we generated 760 segments (380 preterm and 380 term) per training iteration. From each segmented signal, we extracted 57 features per channel as previously described. The corresponding performance metrics are presented in Table IV.
TABLE IV: Classification results for eight models applied to the 2012 TPEHG dataset using fixed 3-minute interval sampling. KLT processed signal variants are evaluated across Accuracy, Precision, Recall, F1 Score, and AUC metrics.
@@ III-C EVALUATION ON THE TPEHG DATASET
These results also highlight the limitations of the available datasets and the challenges posed by sampling strategies. Despite balancing the classes during training, the original imbalance in subject distribution, particularly the low number of preterm cases, can still influence model generalization. This limitation is further emphasized in Section III-D, where we compare our approach to existing models. The analysis reveals that improper oversampling strategies employed in prior works have significantly inflated reported performance, often resulting in overly inflated accuracy metrics that do not accurately reflect real world deployment conditions.
@@
III-D Comparison with existing models
Building upon the methodological concerns raised earlier regarding inflated performance metrics from improper oversampling, our comparative analysis adopts a rigorous approach to benchmark evaluation. We restrict our comparison to studies that have either been methodologically validated by Vandewiele et al. [14] or that explicitly document appropriate sampling strategies in their original publications. Table V presents this curated comparison, including both originally reported results and, where available, the methodologically corrected results from Vandewiele et al. In addition, we incorporate recent classical ML studies that report reproducible methodologies and clinically interpretable outcomes. In addition to the classifiers explored in this study, results from prior works using AdaBoost (AB) are also included for completeness.
@@ III-D Comparison with existing models
TABLE V: Comparison of reported and corrected AUC values from previous studies.
Study Classifier Metric Uncorrected Results Corrected Results
[27, 14] SVM AUC 0.9200 0.6075
[28, 14] AB AUC 0.9860 0.5754
[29, 14] RF AUC 0.9400 0.5231
[30, 14] SVM AUC 0.9900 0.5604
[12, 14] QDA AUC 0.9944 0.6386
[31, 14] RF AUC 0.8880 0.5185
[32] CB AUC NA 0.7000
[33] Bayes AUC NA 0.8400
[17] SVM AUC NA 0.8600
[34] SVM AUC NA 0.8900
[35] QDA AUC NA 0.9030
[36] HVG + SVM AUC NA 0.9700
Proposed CB AUC NA 0.9988
PTB detection is a critical and sensitive clinical task; yet, much of the prior literature has reported overly optimistic results, portraying PTB prediction as a solved problem. In reality, this is far from the truth. As shown in Table V, many earlier studies relied on flawed sampling strategies, particularly improper oversampling and data leakage, leading to dramatically inflated performance metrics. For instance, several reported AUCs exceeding 90% dropped on average by more than 25 percentage points, in some cases falling below 65% when evaluated using proper methodology. This striking discrepancy highlights the extent of overestimation in prior work and underscores the urgent need to reopen and rigorously reevaluate the PTB prediction problem. Notably, the limitations of current datasets continue to be a significant barrier to the development of generalizable and reliable models. This study represents an essential first step in that re-evaluation, while also calling for more extensive datasets and methodological rigor to add credibility to AI based PTB research. In doing so, it aims to build clinical trust by prioritizing explainability, transparency, and interoperability in model design.
@@
IV Conclusion
Preterm birth (PTB) affects approximately 10% of live births and remains the leading cause of death among children under five. Survivors often face lifelong challenges, including developmental delays, chronic diseases, and sensory impairments. Early and accurate prediction of PTB is therefore essential for improving neonatal outcomes and guiding timely interventions. This study proposed a robust and reproducible machine learning (ML) pipeline for PTB prediction using EHG signals, addressing key limitations in prior research, particularly issues with improper oversampling strategies and limited focus on frequency domain features.